OpenStudy (anonymous):

You earn $245 in gross pay a week and 2/5 of your check is taken out for taxes and deductions. How much of your check is taken out for taxes and deductions? A. $49 B. $98 C. $150

OpenStudy (jamesj):

and now you see why math is useful and why you should learn it. so what is 2/5th of $245 ? It is \[\frac{2}{5} \times 245 = \frac{490}{5} = .... what?\]

OpenStudy (anonymous):

i love math i just am not good with these kinds of problem

OpenStudy (jamesj):

so I've shown you how to do it. all you need to do now is the last step and divide 490 by 5.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

You earn $245 a week in gross pay, and 2/5 of your check is taken out for taxes and deductions. How much of your check is taken out for taxes and deductions?